find three consecutive integers such that the sum of the first and the third increased by 8 is 40 more than the second.
Answer by mananth(16946)

let the integers be (n-1) , n, (n+1)
(n-1)+(n+1) +8 =n+40
2n+8=n+40
n=40-8
n=32
The numbers are
31 32 , 33